# Nonconvergence of the Feynman–Dyson diagrammatic perturbation expansion of propagators

### v7

@@ Abstract
Using a general-order ab initio many-body Green’s-function method for molecules, we numerically illustrate several pathological behaviors of the Feynman–Dyson diagrammatic perturbation expansion of one-particle many-body Green’s functions as electron Feynman propagators. (i) The perturbation expansion of the frequency-dependent self-energy is not convergent at the exact self-energy in many frequency domains. (ii) An odd-perturbation-order self-energy has a qualitatively wrong shape and, as a result, many roots of the corresponding Dyson equation are nonphysical in that the poles may be complex or the residues can exceed unity or be negative. (iii) The Dyson equation with an even-order self-energy has roots whose energies approach sheer orbital energy differences of the zeroth-order mean-field theory with little to no electron-correlation effects taken into account. (iv) Infinite partial summation of diagrams by vertex or edge renormalization can exacerbate these problems. Not only do these nonconvergences render higher-order Feynman–Dyson diagrammatic perturbation theory useless for many lower-lying ionization or higher-lying electron-attachment roots, but they also call into question the validity of its combined use with the ansätze requiring the knowledge of all poles and residues. Such ansätze include the Galitskii–Migdal identity, the self-consistent Green’s-function methods, and some models of the algebraic diagrammatic construction.
@@ I Introduction
In an influential paper Dyson (1952), Dyson argued that the Feynman–Dyson diagrammatic perturbation theory for quantum electrodynamics is inherently divergent in the presence of electron-positron pair formations even after mass and charge are renormalized. In another important paper Kohn and Luttinger (1960), Kohn and Luttinger predicted that the finite-temperature diagrammatic perturbation theory for electrons Jha and Hirata (2019); Hirata and Jha (2019, 2020); Jha and Hirata (2020); Hirata (2021a) does not necessarily reduce to the zero-temperature counterpart as the temperature is lowered to zero. In particular, it was argued Kohn and Luttinger (1960); Luttinger and Ward (1960) that the second-order correction to the grand potential can be divergent when the zeroth-order wave function is degenerate, even though the same correction is finite at zero temperature. These predictions have been confirmed both analytically and numerically Hirata (2021b, 2022). Other instances of breakdown of diagrammatic perturbation theory have also been reported Lani et al. (2012); Schäfer et al. (2013); Janiš and Pokorný (2014); Stan et al. (2015); Kozik et al. (2015); Rossi and Werner (2015); Schäfer et al. (2016); Tarantino et al. (2017); Gunnarsson et al. (2017).
@@
In this article, we reveal and analyze several additional pathological behaviors of the Feynman–Dyson diagrammatic perturbation expansions of one-particle many-body Green’s functions or Feynman propagators Feynman (1948); Dyson (1949a, b); Schwinger (1951); Sakurai (1967); March et al. (1967); Mattuck (1992); Dyson (1993); Fetter and Walecka (2003); Kadanoff and Baym (2018). Our analysis is based on the ab initio electron propagators for molecules Linderberg and Öhrn (1965); Hedin (1965); Öhrn and Linderberg (1965); Linderberg and Öhrn (1967); Goscinski and Lukman (1970); Doll and Reinhard (1972); Pickup and Goscinski (1973); Simons and Smith (1973); Yarlagadda et al. (1973); Linderberg and Öhrn (1973); Tsui and Freed (1974); Paldus and Čížek (1975); Cederbaum (1975); Cederbaum and Domcke (1977); Simons (1977); Herman et al. (1978, 1980a, 1980b); Baker and Pickup (1980); Öhrn and Born (1981); Jørgensen and Simons (1981); Schirmer (1982); Schirmer et al. (1983); von Niessen et al. (1984); Prasad et al. (1985); Hybertsen and Louie (1986); Oddershede (1987); Kutzelnigg and Mukherjee (1997); Aryasetiawan and Gunnarsson (1998); Ortiz (1999); Onida et al. (2002); Ortiz (2013); Willow et al. (2013); Hirata et al. (2015, 2017); Johnson et al. (2018); Doran and Hirata (2019); Opoku et al. (2021); Ortiz (2022); Opoku et al. (2022, 2023a, 2023b) as we can take advantage of several independent methods that can determine the poles and residues of their exact (finite-basis-set) Green’s functions Hirata et al. (2000). We also have an algorithm that can evaluate perturbation corrections to their frequency-dependent self-energy and Green’s function at any arbitrary order and frequency Hirata et al. (2017). The conclusions drawn here, however, should be valid for other systems that are studied by the same theory, such as anharmonic molecular vibrations Hermes and Hirata (2013), anharmonic lattice vibrations Qin and Hirata (2020), energy bands in solids Suhai (1983); Sun and Bartlett (1996a, b); Hirata and Bartlett (2000); Hirata and Shimazaki (2009); Willow et al. (2014); Hirata et al. (2023), finite nuclei, and nuclear matter Day (1967); Rowe (1968); Baldo (1999).
@@
Specifically, we show that (i) the perturbation expansion of the frequency-dependent self-energy is nonconvergent at the exact self-energy in many domains of frequency. One notable exception is when the frequency falls in the central domain that encloses zero frequency and most principal roots. (ii) An odd-perturbation-order self-energy has a qualitatively wrong shape except in the central domain. While the diagonal exact self-energy is monotonically decreasing within each frequency bracket separated by its singularities, the diagonal odd-order self-energy can be convex, concave, or monotonically increasing in each bracket. As a result, many roots of the corresponding Dyson equation can be complex and thus nonphysical; when they are real, the corresponding residue can fall outside of the valid range of zero to one. (iii) An even-order self-energy has a qualitatively correct, monotonically decreasing form within each bracket demarcated by its singularities. However, these singularities become more vertical with increasing perturbation order and they occur at orbital energy differences of the zeroth-order mean-field theory. Consequently, the roots of the corresponding Dyson equation approach the sheer orbital energy differences with little to no electron-correlation effects taken into account.
@@
None of these problems is detected in the exact (finite-basis-set) self-energy or Green’s function according to their calculations performed by the full configuration-interaction (FCI) or full equation-of-motion coupled-cluster (EOM-CC) method Hirata et al. (2017, 2000). However, (iv) partial summations of the perturbation corrections up to an infinite order by vertex or edge renormalization tend to exhibit the same pathological behaviors with even greater severity.
@@
Together, these failures pose immediate difficulties when applying higher-than-second-order Feynman–Dyson perturbation theory to the Green’s-function methods that are predicated on the knowledge of all poles and residues. Such methods include the Galitskii–Migdal identity Galitskii and Migdal (1958); Koltun (1972); Holleboom and Snijders (1990); Ortiz (1995), self-consistent Green’s-function methods Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck et al. (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser et al. (2017); Tarantino et al. (2017); Coveney and Tew (2023), and some models of the algebraic diagrammatic construction (ADC) Schirmer (1982); Schirmer et al. (1983); Deleuze et al. (1995); Deleuze (2003). The Luttinger–Ward functional Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik et al. (2015); Rossi and Werner (2015); Gunnarsson et al. (2017); Lin and Lindsey (2018), which serves as a basis of the self-consistent Green’s-function methods and dynamical mean-field theory (DMFT) Georges et al. (1996), may also be negatively impacted by these failures. They also call into question the robustness of the Feynman–Dyson perturbation expansions of Feynman propagators as the mathematical foundation of diagrammatic or perturbative quantum field theory Feynman (1948); Dyson (1949a, b); Schwinger (1951); Sakurai (1967); Dyson (1993).
@@ II.1 Formalisms
An electron propagator is defined in the time ($t$) domain as a time-ordered sum of Green’s functions,
@@
where $\theta(t)$ is the Heaviside step function, $\Psi_{0}$ and $E_{0}$ are the exact wave function and energy for the $N$ electron ground state, and $\hat{p}^{\dagger}$ and $\hat{q}$ are the electron creation and annihilation operators. It describes the probability of an electron (hole) traveling from the $p$th ($q$th) to $q$th ($p$th) spinorbital in time $t$.
@@
where $\eta$ is a positive infinitesimal, $I$ sums over all $N-1$ electron exact states, and $A$ runs over all $N+1$ electron exact states. The first term diverges whenever $\omega$ coincides with an exact ionization potential (IP), whereas the second term has a pole at an exact electron-attachment energy (EA), apart from their signs. The primary utility of the electron propagator for molecules is the direct determination of IPs and EAs for both principal (Koopmans) and satellite (shakeup or non-Koopmans) states. It should be noted, however, the distinction between the principal and satellite roots is not definite, and we use these labels rather loosely in this article.
@@
where $\bm{1}$ and $\bm{\epsilon}$ are, respectively, the unit matrix and the diagonal matrix of $\epsilon_{p}$, which are of the same rank as $\bm{G}(\omega)$ or $\bm{\Sigma}(\omega)$. Hence, one can determine all poles and residues of $\bm{G}(\omega)$ by solving
@@
In addition to IPs and EAs, the exact total energy is gleaned from $\bm{G}(\omega)$ by using the Galitskii–Migdal identity Galitskii and Migdal (1958); Koltun (1972); Holleboom and Snijders (1990); Ortiz (1995),
@@
In the diagonal approximation Hirata et al. (2017) to the self-energy, the inverse Dyson equation now simplifies to
@@ II.2 Numerical results
Figure 2: The exact diagonal $G_{33}(\omega)$ as a function of $\omega$ for the BH molecule (1.232 Å in the minimal basis set; the third spatial orbital in the increasing order of orbital energy corresponds to the HOMO). The exact IPs and EAs obtained by FCI are superposed (as open circles), occurring at the poles of the Green’s function. The HF and FCI energies are $-24.752788\,E_{\text{h}}$ and $-24.809940\,E_{\text{h}}$, respectively.
@@
In Fig. 2 is plotted the third diagonal element of the exact $\bm{G}(\omega)$ matrix as functions of $\omega$ for the boron hydride molecule with the bond length of 1.232 Å in the minimal basis set. The third element corresponds to the highest-occupied molecular orbital (HOMO). Other elements of $\bm{G}(\omega)$ are omitted to avoid clutter. The exact $\bm{G}(\omega)$ was obtained by literally evaluating Eq. (2) using a determinant-based FCI program Hirata et al. (2017, 2015). The figure confirms the well-known fact Walter et al. (1984) that the function is divided by singularities into consecutive regions or brackets, within each of which it is a monotonically decreasing function of $\omega$.
@@
In the same figure are superposed the exact IPs and EAs (signs reversed; 300 each) obtained by the determinant-based FCI program Hirata et al. (2000). They coincide with the poles of $G_{33}(\omega)$ as they should. There are some IPs and EAs that apparently lack matching poles, but they correspond either to the poles of other elements of $\bm{G}(\omega)$ (e.g., for point-group symmetry reasons) or to nearly vertical poles (with nearly zero residues) that have fallen through the $\omega$ mesh used for plotting Fig. 2. These are merely some minor issues associated with the algorithms adopted, and there is no formal or serious practical problem in the exact MBGF method.
@@
Figure 3: Eigenvalues of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ as a function of $\omega$ for the BH molecule. The exact IPs and EAs obtained by FCI are superposed (as open circles), and they coincide with the roots of the inverse Dyson equation (the intersections of the eigenvalues with the diagonal line).
@@
Figure 3 plots eigenvalues of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ matrix as functions of $\omega$. Of the six eigenvalues the lowest one is not visible in this plot and the second and third highest ones are degenerate. Each root of the inverse Dyson equation [Eq. (9)] is expected at an intersection of the eigenvalues of $\bm{\epsilon}+\bm{\Sigma}(\omega)$ and the diagonal line $\omega$. In fact, the IPs and EAs obtained from FCI are seen to occur precisely at these intersections. The few IPs and EAs appearing to occur away from any intersection are likely due to the nearly vertical poles with nearly zero residues, which are thus undetected by an $\omega$ mesh. This issue may be viewed as a weakness of the graphical method Walter et al. (1984) of solving the inverse Dyson equation if one is concerned with determining all the roots, but may also be considered an advantage if one prefers to avoid numerous “phantom” roots with zero or near zero residues.
@@
Figure 4 is a histogram of the poles of exact $\bm{G}(\omega)$; the height of each impulse is the corresponding residue [Eq. (10)]. There are numerous poles outside of this graph (see Sec. IV). Generally, the poles and residues can be determined by the graphical or arrow-matrix diagonalization method Walter et al. (1984). We used neither; the poles $\{\omega_{q}\}$ in their entirety were obtained by the FCI method Hirata et al. (2000) and verified by substitution to the inverse Dyson equation [Eq. (9)]. The residue at $\omega=\omega_{q}$ was then computed by evaluating Eq. (10). The derivative of $\bm{\Sigma}$ with respect to $\omega$ can be taken analytically by
@@
However, the last expression is divergent at every pole $\omega=\omega_{q}$. We therefore approximated this derivative as an average of the derivatives at $\omega=\omega_{q}\pm 10^{-9}\,E_{\text{h}}$.
The residues thus obtained correctly fall in the range of zero to one. They also satisfy the particle number sum rule [Eq. (11)] and Galitskii–Migdal identity [Eq. (12)] with the precision of $10^{-7}$ and $10^{-6}\,E_{\text{h}}$, respectively, which is a numerical manifestation of the fact that the exact Green’s function obeys the Baym–Kadanoff conservation laws Baym and Kadanoff (1961); Baym (1962); Kadanoff and Baym (2018); Dahlen and van Leeuwen (2005). Very many tiny contributions from satellite roots are crucial for these identities to be accurately satisfied.
Overall, the exact finite-basis-set Green’s function and self-energy are well behaved, satisfying conservation laws and yielding results that are in exact numerical agreement with alternative methods such as FCI or EOM-CC. Therefore, the pathological behaviors we are about to discuss are exclusively ascribed to perturbation theory.

### v8

@@ Abstract
Using a general-order ab initio many-body Green’s function method, we numerically illustrate several pathological behaviors of the Feynman–Dyson diagrammatic perturbation expansion of one-particle many-body Green’s functions as electron Feynman propagators. (i) The perturbation expansion of the frequency-dependent self-energy is not convergent at the exact self-energy in many frequency domains. (ii) An odd-perturbation-order self-energy has a qualitatively wrong shape and, as a result, many roots of the corresponding Dyson equation are nonphysical in that the poles may be complex or residues can exceed unity or be negative. (iii) A higher even-order self-energy consists of vertical lines at many frequencies, predicting numerous phantom poles with zero residues. (iv) Infinite partial resummations of diagrams by vertex or edge renormalization tend to exacerbate these pathologies. (v) The nonconvergence is caused by the nonanalyticity of the rational-function form of the exact Green’s function at many frequencies, where the radius of convergence of its Taylor expansion is zero. This is consistent with the fact that (vi) Padé approximants (power-series expansions of a rational function) can largely restore the correct shape and poles of the Green’s function. Nevertheless, not only does the nonconvergence render higher-order Feynman–Dyson diagrammatic perturbation theory useless for many lower-lying ionization or higher-lying electron-attachment states, but it also calls into question the validity of its combined use with the ansätze requiring the knowledge of all poles and residues. Such ansätze include the Galitskii–Migdal identity, the self-consistent Green’s function methods, and some models of the algebraic diagrammatic construction.
@@ I Introduction
In an influential paper Dyson (1952), Dyson argued that the Feynman–Dyson diagrammatic perturbation theory for quantum electrodynamics is inherently divergent in the presence of electron-positron pair formations even after mass and charge are renormalized. In another important paper Kohn and Luttinger (1960), Kohn and Luttinger predicted that the finite-temperature diagrammatic perturbation theory for electrons Jha and Hirata (2019); Hirata and Jha (2019, 2020); Jha and Hirata (2020); Hirata (2021a) does not necessarily reduce to the zero-temperature counterpart as the temperature is lowered to zero. In particular, it was argued Kohn and Luttinger (1960); Luttinger and Ward (1960) that the second-order correction to the grand potential can be divergent when the zeroth-order wave function is degenerate, whereas the same correction is finite at zero temperature. These predictions have been confirmed both analytically and numerically Hirata (2021b, 2022). Other instances of breakdown of diagrammatic perturbation theory have also been reported Lani et al. (2012); Schäfer et al. (2013); Janiš and Pokorný (2014); Stan et al. (2015); Kozik et al. (2015); Rossi and Werner (2015); Schäfer et al. (2016); Tarantino et al. (2017); Gunnarsson et al. (2017).
@@
In this article, we reveal and analyze several additional pathological behaviors of the Feynman–Dyson diagrammatic perturbation expansions of one-particle many-body Green’s functions (MBGF) or Feynman propagators Feynman (1948); Dyson (1949a, b); Schwinger (1951); Sakurai (1967); March et al. (1967); Mattuck (1992); Dyson (1993); Fetter and Walecka (2003); Kadanoff and Baym (2018). Our analysis is based on the ab initio electron propagators for molecules Linderberg and Öhrn (1965); Hedin (1965); Öhrn and Linderberg (1965); Linderberg and Öhrn (1967); Goscinski and Lukman (1970); Doll and Reinhard (1972); Pickup and Goscinski (1973); Simons and Smith (1973); Yarlagadda et al. (1973); Linderberg and Öhrn (1973); Tsui and Freed (1974); Paldus and Čížek (1975); Cederbaum (1975); Cederbaum and Domcke (1977); Simons (1977); Herman et al. (1978, 1980a, 1980b); Baker and Pickup (1980); Öhrn and Born (1981); Jørgensen and Simons (1981); Schirmer (1982); Schirmer et al. (1983); von Niessen et al. (1984); Prasad et al. (1985); Hybertsen and Louie (1986); Oddershede (1987); Kutzelnigg and Mukherjee (1997); Aryasetiawan and Gunnarsson (1998); Ortiz (1999); Onida et al. (2002); Ortiz (2013); Willow et al. (2013); Hirata et al. (2015, 2017); Johnson et al. (2018); Doran and Hirata (2019); Opoku et al. (2021); Ortiz (2022); Opoku et al. (2022, 2023a, 2023b) as we can take advantage of several independent methods that can determine the poles and residues of their exact (finite-basis-set) Green’s functions Hirata et al. (2000) as well as the algorithm that can evaluate perturbation corrections to their frequency-dependent self-energy and Green’s function at any arbitrary order and frequency Hirata et al. (2017). The conclusions drawn here, however, should be valid for other systems that are studied by the same theory, such as anharmonic molecular vibrations Hermes and Hirata (2013), anharmonic lattice vibrations Qin and Hirata (2020), energy bands in solids Suhai (1983); Sun and Bartlett (1996a, b); Hirata and Bartlett (2000a); Hirata and Shimazaki (2009); Willow et al. (2014); Hirata et al. (2023), finite nuclei, and nuclear matter Day (1967); Rowe (1968); Baldo (1999).
@@
Specifically, we show that (i) the perturbation expansion of the frequency-dependent self-energy is nonconvergent at the exact self-energy in many domains of frequency. (ii) An odd-perturbation-order self-energy has a qualitatively wrong shape except in the central domain that encloses zero frequency and most principal roots. While the diagonal exact self-energy is monotonically decreasing within each frequency bracket separated by its singularities, the diagonal odd-order self-energy can be convex, concave, or monotonically increasing in each bracket. As a result, many roots of the corresponding Dyson equation can be complex and thus nonphysical; when they are real, the corresponding residue can fall outside of the valid range of zero to one. (iii) An even-order self-energy may have a qualitatively correct, monotonically decreasing form within each bracket demarcated by its singularities. However, they become more and more vertical with increasing perturbation order, predicting numerous “phantom” poles with zero residues.
@@
None of these problems is detected in the exact self-energy or Green’s function, which can be determined (within a finite basis set) by the full configuration-interaction (FCI) or, equivalently, by the full equation-of-motion coupled-cluster (EOM-CC) method Hirata et al. (2017, 2000), which may be viewed as an infinite resummation of the Feynman–Dyson perturbation corrections Hirata et al. (2015). However, (iv) other infinite partial resummations of propagator diagrams by vertex or edge renormalization such as the Tamm–Dancoff approximation Day (1967); Kadanoff and Baym (2018); Linderberg and Öhrn (1967); Purvis and Öhrn (1974); Schirmer and Cederbaum (1978); Walter and Schirmer (1981); Baldo (1999) and self-consistent second-order Green’s function method Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck et al. (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser et al. (2017); Tarantino et al. (2017); Coveney and Tew (2023) are shown to exhibit the same pathologies with even greater severity.
We also elucidate (v) the cause of the nonconvergence to be the fact that the definition of the Green’s function is nonanalytic at many frequencies, where the radius of convergence of its Taylor expansions is zero. In other words, the breakdown of the Feynman–Dyson perturbation theory occurs due to a purely mathematical reason, i.e., that the rational-function form of the exact Green’s function does not lend itself to a converging Taylor expansion, even though physical information such as the converging perturbation expansions of all state energies are actually contained in it. Consequently, (vi) the general mathematical technique of Padé approximants Goscinski (1967); Goscinski and Lukman (1970); Brändas and Goscinski (1970); Linderberg and Ratner (1970); Goscinski and Brändas (1971); Bartlett and Brändas (1972, 1973); Laidig et al. (1985); Bender and Orszag (1999); Hirata and Bartlett (2000b), which are power-series expansions of a rational function, can extract this physical information, systematically and rapidly restoring the correct shape and poles of the Green’s function.
@@
Nevertheless, the nonconvergence of the Feynman–Dyson perturbation theory, despite being the mathematical foundation of quantum field theory Feynman (1948); Dyson (1949a, b); Schwinger (1951); Sakurai (1967); Dyson (1993), poses difficulties when applying higher-than-second-order approximations to the Green’s function methods that are predicated on the knowledge of all poles and residues. Such methods include the Galitskii–Migdal identity Galitskii and Migdal (1958); Koltun (1972); Holleboom and Snijders (1990); Ortiz (1995), self-consistent Green’s function methods Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck et al. (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser et al. (2017); Tarantino et al. (2017); Coveney and Tew (2023), and some models of the algebraic diagrammatic construction (ADC) Schirmer (1982); Schirmer et al. (1983); Deleuze et al. (1995); Deleuze (2003). The Luttinger–Ward functional Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik et al. (2015); Rossi and Werner (2015); Gunnarsson et al. (2017); Lin and Lindsey (2018), which serves as a basis of the self-consistent Green’s function methods and dynamical mean-field theory (DMFT) Georges et al. (1996), may also be negatively impacted by these failures.
@@ II.1 Formalisms
An electron Feynman propagator is defined in the time ($t$) domain as a time-ordered sum of Green’s functions,
@@
where $\theta(t)$ is the Heaviside step function, $\Psi_{0}$ and $E_{0}$ are the exact wave function and energy for the $N$ electron ground state, and $\hat{p}^{\dagger}$ and $\hat{q}$ are the electron creation and annihilation operators. It describes the probability of an electron (hole) traveling from the $p$th ($q$th) to $q$th ($p$th) spinorbital in time $t$ ($-t$).
@@
where $\eta$ is a positive infinitesimal, $I$ sums over all $N-1$ electron exact states, and $A$ runs over all $N+1$ electron exact states. (Here, $\eta$ is just mathematical convenience and has nothing to do with a lifetime of spectral bandwidth, which are infinite and zero, respectively. Nonetheless, the present formulation can be applied to solids without any modification, in which a spectral bandwidth emerges naturally and rigorously as a manifold of closely packed $\delta$-function peaks.)
The first term diverges whenever $\omega$ coincides with an exact ionization potential (IP), whereas the second term has a pole at an exact electron-attachment energy (EA), apart from their signs. The primary utility of the electron propagator for molecules and solids is the direct determination of IPs and EAs (or electron-correlated energy bands) for both principal (Koopmans) and satellite (shakeup or non-Koopmans) states. It should be noted, however, the distinction between the principal and satellite roots is not definite, and we use these labels rather loosely in this article.
@@
where $\bm{1}$ and $\bm{\epsilon}$ are, respectively, the unit matrix and the diagonal matrix of $\epsilon_{p}$, which are of the same rank as $\bm{G}(\omega)$ or $\bm{\Sigma}(\omega)$.
One can therefore determine all poles and residues of $\bm{G}(\omega)$ by solving
@@
In addition to IPs and EAs, the exact total energy is gleaned from $\bm{G}(\omega)$ with the aid of the Galitskii–Migdal identity Galitskii and Migdal (1958); Koltun (1972); Holleboom and Snijders (1990); Ortiz (1995),
@@
In the diagonal approximation Hirata et al. (2017) to the self-energy, the inverse Dyson equation simplifies to
@@ II.2 Numerical results
Figure 2: The third diagonal element of the exact $\bm{G}(\omega)$ as a function of $\omega$ for the BH molecule (1.232 Å in the minimal basis set; the third spatial orbital in the increasing order of orbital energy corresponds to the HOMO). The exact IPs and EAs obtained by FCI are superposed (as open circles), occurring at the poles of the Green’s function. The HF and FCI energies are $-24.752788\,E_{\text{h}}$ and $-24.809940\,E_{\text{h}}$, respectively.
@@
In Fig. 2 is plotted the third diagonal element of the exact $\bm{G}(\omega)$ matrix as functions of $\omega$ for the boron hydride molecule with the bond length of 1.232 Å in the minimal basis set. The third element corresponds to the highest-occupied molecular orbital (HOMO). The other elements of $\bm{G}(\omega)$ are omitted to avoid clutter. The exact $\bm{G}(\omega)$ was obtained by literally evaluating Eq. (2) using a determinant-based FCI program Hirata et al. (2017, 2015). The figure confirms the well-known fact Walter et al. (1984) that the function is divided by singularities into consecutive regions or brackets, within each of which it is a monotonically decreasing, “shoulder-like” function of $\omega$.
@@
In the same figure are superposed the exact IPs and EAs (signs reversed; 300 each) obtained by the determinant-based FCI program Hirata et al. (2000). They coincide with the poles of $G_{33}(\omega)$ as they should. There are some IPs and EAs that apparently lack matching poles, but they correspond either to the poles of other elements of $\bm{G}(\omega)$ or to nearly vertical poles (with nearly zero residues) that have fallen through the $\omega$ mesh used for plotting Fig. 2. These are merely some minor issues associated with the algorithms adopted, and there is no pathological behavior detectable in the exact MBGF, which reproduces the results of FCI.
@@
Figure 3: Eigenvalues of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ as a function of $\omega$ for the BH molecule. The exact IPs and EAs obtained by FCI are superposed (as open circles), and they coincide with the roots of the inverse Dyson equation, occurring at the intersections of the eigenvalues with the diagonal $\omega$ line; see Eq. (9).
@@
Figure 3 plots eigenvalues of the exact $\bm{\epsilon}+\bm{\Sigma}(\omega)$ matrix as functions of $\omega$. Of the six eigenvalues the lowest one is not visible in this plot and the second and third highest ones are degenerate. Each root of the inverse Dyson equation is expected at an intersection of the eigenvalues of $\bm{\epsilon}+\bm{\Sigma}(\omega)$ and the diagonal line $\omega$ as per Eq. (9). In fact, the IPs and EAs obtained from FCI are seen to occur precisely at these intersections. The few IPs and EAs appearing to occur away from any intersection are likely due to the nearly vertical poles with nearly zero residues, which are thus undetected by an $\omega$ mesh. This issue may be a weakness of the graphical method Walter et al. (1984); Piecuch et al. (1990) of solving the inverse Dyson equation if one is concerned with determining all roots, but it by no means signals any fundamental problem in the theory.
@@
Figure 4 is a histogram of the poles of the exact $\bm{G}(\omega)$; the height of each impulse is the corresponding residue [Eq. (10)]. There are numerous poles outside of this graph (see Sec. IV). Generally, the poles and residues can be determined by the graphical Walter et al. (1984); Piecuch et al. (1990) or arrow-matrix diagonalization method Walter et al. (1984). In this study, we used neither; the poles $\{\omega_{q}\}$ in their entirety were first obtained by FCI Hirata et al. (2000) and verified by substitution to the inverse Dyson equation [Eq. (9)]. The residue at $\omega=\omega_{q}$ was then computed by evaluating Eq. (10). The derivative of $\bm{\Sigma}$ with respect to $\omega$ can be taken analytically by
@@
However, the last expression is ill-conditioned at every pole $\omega=\omega_{q}$. We therefore approximated this derivative as an average of the derivatives at $\omega=\omega_{q}\pm 10^{-9}\,E_{\text{h}}$.
The residues thus obtained correctly fall in the range of zero to one for the BH molecule. They also satisfy the particle number sum rule [Eq. (11)] and Galitskii–Migdal identity [Eq. (12)] with the precision of $10^{-7}$ and $10^{-6}\,E_{\text{h}}$, respectively, which is a numerical manifestation of the fact that the exact Green’s function obeys the Baym–Kadanoff conservation laws Baym and Kadanoff (1961); Baym (1962); Kadanoff and Baym (2018); Dahlen and van Leeuwen (2005). Very many tiny contributions from satellite roots are crucial for these identities to be accurately satisfied, and including only principal roots results in severe errors even in this tiny system.
Overall, the exact finite-basis-set Green’s function and self-energy are well-behaved, satisfying conservation laws and yielding results that are in exact numerical agreement with alternative methods such as FCI or full EOM-CC. Therefore, the pathological behaviors we are about to discuss are exclusively ascribed to their perturbation expansions.